Question 6 1 pts 8 years ago you purchased ABC stock for $1 per share. Today, you sell the stock for $1 per share. What is your average annual return on this investment? Answer in decimals (not %) and round to 4 decimals. For example, 0.1234.
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

1=1*(1+r/100)^8

(1+r/100)=(1)^(1/8)

(1+r/100)=1

r=1-1

=0